OpenStudy (anonymous):

-10x+300=0 find x and y intercept

OpenStudy (timmyk14):

The y-intercept is 300, and to find the value of x you solve for x. -10x + 300 = 0 - 300 -300 -10x = -300 Divide both sides by -10 x = 30 Now to check our work: -10(-30) +300 = 0 -300 + 300 = 0

OpenStudy (timmyk14):

I know this is a bit late but the 30 is not supposed to be a negative sorry. It should look like this: -10(30) + 300 = 0 -300 + 300 = 0 0 = 0